Greenhab Report – April 30th

Crew 211 GreenHab Report – 30-APR-2019

GreenHab Officer: Chloé Peduzzi

Environmental control: Ambient with door open (but we had to close the door when there was rain and to much wind) + fan turned up to 3

Shade cloth: 80%

Average temperatures:

Low: 17°C

High: 35°C

Hours of supplemental light: 0

Daily water usage for crops: 3 gallons + 2 gallons

Water in Blue Tank: 75 gallons

Time(s) of watering for crops: 10:00 + 19:00

Changes to crops: All plants look healthy. Nevertheless the tomato plant of the GreenHab still seems a little stressed (curved leaf), I will water it accordingly over the next couple of days.

Narrative: It was also cooler today, the rain and the wind have helped to decrease the temperature. We have harvested Basil and Tomatoes (from plants brought by our crew) for the dinner.

Harvest: 13.8g Basil + 105g Tomatoes (from plants brought by our crew)

Support/supplies needed: I won’t have enough water for 1 week long, as it was really hot this past week, and I had to water plants accordingly.

Astronomy Report – April 30th

Astronomy Report

Name: Lieffrig Eléonore

Crew: 211
Date: 30 April 2019


Solar Features Observed: Granules, great activity on the surface of the sun but no prominences observed.

Problems Encountered: The tracking system doesn’t work despite the new procedure given by Mr. Detterline. The tracking was set on sideral, therefore we changed it to solar, but the telescope didn’t even move.

Sol Summary – April 30th

Crew 211 Sol 9 Summary Report 30-APR-2019

Sol: 9

Summary Title: Storm over Mars

Author’s name: Carl-Henrik Dahlqvist

Mission Status: Nominal

Sol Activity Summary: As we had numerous showers during the night, the morning EVA was unfortunately postponed to the afternoon and then cancelled as the weather deteriorated in the afternoon. We therefore worked all day on our experiments. We had an interesting meeting after the lunch to discuss with the crew journalist some articles of our Martian constitution. The experiment of the Commander and the Vice-Commander are now almost ready for field experiment.

Look Ahead Plan: Depending on the weather, we will try again to do a morning EVA to Kissing Camel and an afternoon EVA to Candor Chasma, as the end of our stay is approaching. We will in parallel work on our respective experiments and hopefully be able to start in the afternoon the field experiments with the cubelanders and the geolocation antennas.

Anomalies in work: No issues.

Weather: Showers, mostly cloudy and very windy

Crew Physical Status: Good

EVA: The EVA has been cancelled due to bad weather.

Reports to be filed: Journalist Report, GreenHab Report, Operations Report, Science Report, EVA Request, Daily Photo Report

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site